# mywords.dic

### v1

@@ A.3.3 GSPO
GSPO discards more than 10% of responses with non-zero advantages, resulting in high data waste and training efficiency.
@@ A.5 Hypothetical Reasoning Process for Dynamic-Adaptive Clipping Bounds
For this clipping ratio to be mathematically valid and meaningful, we must adhere to the fundamental constraints of probability distributions and their ratios. Specifically, as shown in Equation 23, the probabilities $p(x)$ and $q(x)$ must be nonnegative and less than or equal to 1. Consequently, the probability ratio $r(x)$ must also be non-negative, and the clipping hyper-parameter $\epsilon$ must be non-negative.
@@ A.8 The Template for Experiments
A.9 Training Common Setting
@@ A.10 The detailed Performance of Models
Figure 5 shows the performance of the model using the average Avg@1 metric. This metric is calculated by averaging the greedy decoding (Avg@1) results on the MATH500, AMC23, AIME24, and AIME25 benchmarks and represents the best performance of the model. As can be seen in the figure, DCPO consistently outperformed GRPO and DAPO across all models throughout the experiment. Furthermore, its overall performance was more stable than the baseline, without the significant performance drop seen in GRPO or the significant fluctuations seen in DAPO. This suggests that DCPO can stimulate stronger model reasoning capabilities during training, providing greater potential for the model to explore even better performance.
@@
Figure 6 shows the Avg@32 performance of DCPO with GRPO and DAPO at different model sizes on AMC23, AIME24, and AIME25 benchmarks, clearly demonstrating DCPO’s exceptional robustness and stability. In contrast, GRPO exhibits significant fluctuation and occasional performance collapse. By GRPO, the Qwen2.5-14B model even experiences a significant performance drop after approximately 170 training steps. Across all evaluated models (from 1.5B to 14B), the DCPO performance trajectory exhibits a consistently upward trend with minimal fluctuation. Even in the early stages of training, DCPO maintains the fastest growth rate and maintains its upward trend. Furthermore, compared to DAPO, DCPO achieves superior performance while saving half of training time and resources. These experimental results demonstrate the effectiveness and robustness of dynamic-adaptive clipping and smoothing advantage standardization to improve the exploration at the response level and the label level of diversity.
@@
The behavior of DAPO and DCPO is model dependent. In the Instruct model (Qwen2.5-Math-1.5B), DAPO, DCPO, and GRPO follow similar declining trajectories. In contrast, for the larger base models (Qwen2.5-3B, Qwen2.5-Math-7B, and Qwen2.5-14B), the trends differ significantly. DAPO maintains the highest entropy value, showing a clear upward trend accompanied by large fluctuations. In contrast, DCPO reaches a moderate entropy level: its curve is significantly smoother and remains confined to a bounded band between the low-entropy region of GRPO and the high-variance region of DAPO.

### v2

@@ A.3.3 GSPO
GSPO discards more than 10% of responses with non-zero advantages, resulting in high data waste and training efficiency. Although GRPO discards responses with sequence-level high variance, it still retains a large number of responses with token-level high entropy, resulting in training instability.
@@ A.5 Hypothetical Reasoning Process for Dynamic-Adaptive Clipping Bounds
For this clipping ratio to be mathematically valid and meaningful, we must adhere to the fundamental constraints of probability distributions and their ratios. Specifically, as shown in Equation 23, the probabilities $p(x)$ and $q(x)$ must be non-negative and less than or equal to 1. Consequently, the probability ratio $r(x)$ must also be non-negative, and the clipping hyper-parameter $\epsilon$ must be non-negative.
@@ A.8 The Template for Experiments
A.9 Training Setting
We set a symmetric clipping threshold of $\epsilon=0.2$ in GRPO and lower/upper clipping thresholds $(0.2,0.28)$ with an additional 512-token soft-punish cache and the maximum probability ratio 10 for negative advantage for DAPO, lower/upper clipping thresholds (3e4,4e4)(3e-4,4e-4) for GSPO, and low/high clipping thresholds $(0.16,0.2)$ for DCPO, preserving the same clipping points $(q(x),r(x))=(\frac{1}{1+\epsilon},1+\epsilon)$ and $(q(x),r(x))=(1,1-\epsilon)$ as GRPO (Figure 4(d)), and set the maximum probability ratio to 10 for both positive and negative advantages, different from DAPO.
@@ A.10 The detailed Performance of Models
Figure 5 shows the performance of the model using the average Avg@1 metric. This metric is calculated by averaging the greedy decoding (Avg@1) results on the MATH500, AMC23, AIME24, and AIME25 benchmarks and represents the best performance of the model. As can be seen in the figure, DCPO consistently outperformed GRPO and DAPO across all four models, and outperformed GSPO on the two models where GSPO was evaluated. Furthermore, its overall performance was more stable than the baseline, without the significant performance drop seen in GRPO or the significant fluctuations seen in DAPO. This suggests that DCPO can stimulate stronger model reasoning capabilities during training, providing greater potential for the model to explore even better performance.
@@
Figure 6 shows the Avg@32 performance of DCPO with GRPO, DAPO and (where available) GSPO at different model sizes on AMC23, AIME24, and AIME25 benchmarks, clearly demonstrating DCPO’s exceptional robustness and stability. In contrast, GRPO exhibits significant fluctuation and occasional performance collapse. By GRPO, the Qwen2.5-14B model even experiences a significant performance drop after approximately 170 training steps. Across all evaluated models (from 1.5B to 14B), the DCPO performance trajectory exhibits a consistently upward trend with minimal fluctuation. Even in the early stages of training, DCPO maintains the fastest growth rate and maintains its upward trend. Furthermore, compared to DAPO, DCPO achieves superior performance while saving half of training time and resources. These experimental results demonstrate the effectiveness and robustness of dynamic-adaptive clipping and smoothing advantage standardization to improve the exploration at the response level and the label level of diversity.
@@
The behavior of DAPO and DCPO is model-dependent. In the Instruct model (Qwen2.5-Math-1.5B), DAPO, DCPO, and GRPO follow similar declining trajectories. In contrast, for the larger base models (Qwen2.5-3B, Qwen2.5-Math-7B, and Qwen2.5-14B), the trends differ significantly. DAPO maintains the highest entropy value, showing a clear upward trend accompanied by large fluctuations. On Qwen2.5-Math-7B, the entropy value of GSPO is similar to that of DCPO, and the average metric performance is also comparable. However, on Qwen2.5-14B, the performance of GSPO exhibits a fluctuating downward trend, resulting in poor performance. In contrast, DCPO reaches a moderate entropy level: its curve is significantly smoother and remains confined to a bounded band between the low-entropy region of GRPO and the high-variance region of DAPO.